OpenStudy (anonymous):

Can anyone help me solve this problem ? I have no idea how to do it. An investor invested a total of $1,900 in two mutual funds, one fund earned a 7% profit, while the other earned a 4% profit. If the investor's total profit was $97, how much was invested into each mutual fund?

OpenStudy (allank):

In this case I think we can safely assume that the interest is compounded continuously for a period of time. Assume that the amount of money invested in the 7% fund be x. Thus the total amount of money after some time t is given by the equations: For the 7% A(t) = 1.07t*x For the 4%: A(t) = 1.04t*(1900-x) So the total amount is given by: (1.07t*x)+(1.04t*(1900-x)) The total profit equation is therefore: 97 = 1900-[(1.07t*x)+(1.04t*(1900-x))] Solving for x should give the solution.

OpenStudy (saifoo.khan):

x+y = 1900 -eq1 0.04x + 0.07y = 97 -eq2 Now solve both of them.